OpenStudy (anonymous):

0.2x + 0.4y = 1.7 and 8.3x - 6.3y= -4.3 how to solve using substitution or elimination and then get the y and x intercept for each one ?

OpenStudy (anonymous):

step one is get rid of annoying decimals my multiplying by ten to get \[2x+4y=17\] \[83x-63y=-43\]

OpenStudy (anonymous):

that helps a lot because the decimals are what were giving me a hard time. i am not sure why i have never thought to do that

OpenStudy (anonymous):

should be easy now yes?

OpenStudy (anonymous):

now would i multiply the first equation by 15.75 to eliminate the y variables ?

OpenStudy (anonymous):

we got rid of the decimals, so forget about them

OpenStudy (anonymous):

Wonder where 15.75 comes from..

OpenStudy (anonymous):

this problem is ugly ugly ugly. your math teacher must hate you

OpenStudy (anonymous):

i guess \[15.75\times 4=63\] maybe

OpenStudy (anonymous):

I thi)nk ur right, his math teacher hates him...:-

OpenStudy (anonymous):

so that would clear the y's but make a huge mess with the calculation. there is nothing nice here. everything is relatively prime, so you just have to multiply it all out. i got \[x=\frac{899}{458}\] \[y=\frac{1497}{458}\] but i cheated.

OpenStudy (anonymous):

if i did this by hand i would have made a mistake no doubt about it.

OpenStudy (anonymous):

Its an online college she does hate us. I got 15.75 to elminate -63y. when i do figure out the solution how should i go about getting the y and x intercepts ?

OpenStudy (anonymous):

If u get full marks for the right answer, feed the equations into Wolfram. What do u mean by the x and y intercepts for each one.._

OpenStudy (anonymous):

I need the x and y intercepts for each of the initial equations after i find the solution set

OpenStudy (anonymous):

If u just mean where the lines cut the x and y axes, that-s easy enough: Just set x= 0 to get y and vice versa for each line.

OpenStudy (anonymous):

thanks! i will keep trying with the problem